OpenStudy (anonymous):

find the extreme values of the function and where they occur. x^3-3x^2+1

OpenStudy (mathmath333):

u have to differentiate the function here

OpenStudy (mathmath333):

after differentiating u have to equate it to zero \(\large\tt \begin{align} \color{black}{x^3-3x^2+1 \\~\\ \dfrac{d}{dx}(x^3)-\dfrac{d}{dx}(3x^2)+\dfrac{d}{dx}(1 )=0\\~\\ 3x^2-6x+0=0\\~\\ 3x^2-6x=0\\~\\~\\~\\ \Large \rm x=0\\~\\ or\\~\\ \Large \rm x=2 }\end{align}\)

OpenStudy (mathmath333):

so when x=0 \(\large\tt \begin{align} \color{black}{y=x^3-3x^2+1 \\~\\ y=0^3-3\times 0^2+1 \\~\\ y=1 }\end{align}\) when x=2 \(\large\tt \begin{align} \color{black}{y=x^3-3x^2+1 \\~\\ y=2^3-3\times 2^2+1 \\~\\ y=-3 }\end{align}\) so the extreme points are \(\Large (x,y)=(0,1)(max)\) and \(\Large (x,y)=(2,-3)(min)\)